# Generalizable, Fast, and Accurate DeepQSPR with fastprop

### v1

@@ Introduction
For this unique dataset, execution time trends are inverted. fastprop takes 2 minutes and 44 seconds, of which 1 minute and 44 seconds were spent calculating descriptors for these unusually large molecules. Chemprop completes in 1 minute and 17 seconds, on par with the training time of fastprop without considering descriptor calculation.
Classification Datasets
See Table LABEL:classification_results_table for a summary of all the classification dataset results. Especially noteworthy is the performance on QuantumScents dataset, which outperforms the best literature result. Citations for the datasets themselves are included in the sub-sections of this section.
Benchmark | Samples (k) | Metric | Literature Best | fastprop | Chemprop |
HIV | ~41 | AUROC | 0.81a | 0.81 | 0.77a |
QuantumScents | ~3.5 | AUROC | 0.88b | 0.91 | 0.85b |
SIDER | ~1.4 | AUROC | 0.67c | 0.64 | 0.65c |
Pgp | ~1.3 | AUROC | 0.94e | 0.92 | 0.89e |
ARA | ~0.8 | Accuracy | 91d | 89 | 82* |
a [9] b [36] c [8] d [37] e [38] * These results were generated for this study.
HIV Inhibition
Originally compiled by Riesen and Bunke [39], this dataset includes the reported HIV activity for approximately 41k small molecules. This is an established benchmark in the molecular property prediction community and the exact version used is that which was standardized in MoleculeNet [24]. This dataset is unique in that the labels in the original study include three possible classes (a multiclass) regression problem whereas the most common reported metric is instead lumping positive and semi-positive labels into a single class to reduce the task to binary classification; both are reported here. UniMol is again used as a point of comparison, and thus an 80/10/10 scaffold-based split with three repetitions is used.
For binary classification fastprop’s AUROC of 0.81 0.04 matches the literature best UniMol with and 0.808 0.003 [9]. This corresponds to an accuracy of 96.8+/-1.0% for fastprop, which taken in combination with AUROC hints that the model is prone to false positives. Chemprop performs worse than both of these models with a reported AUROC of 0.771 0.005.
When attempting multiclass classification, fastprop maintains a similar AUROC of 0.818+/-0.019 AUROC Accuracy suffers a prodigious drop to 42.8 7.6%, now suggesting that the model is prone to false negatives. Other leading performers do not report performance metrics on this variation of the dataset.
QuantumScents
@@
In the reference study, Chemprop achieved an AUROC of 0.85 with modest hyperparameter optimization and an improved AUROC of 0.88 by incorporating the atomic descriptors calculated as part of QuantumScents. fastprop, using neither these descriptors nor the 3D structures, outperforms both models with an AUROC of 0.910+/-0.001 with only descriptors calculated from the molecules’ SMILES. The GitHub repository contains an example of generating custom descriptors incorporating the 3D information from QuantumScents and passing these to fastprop; impact on the performance was negligible.
SIDER
First described by Kuhn et al. in 2015 [41], the Side Effect Resource (SIDER) database has become a standard property prediction benchmark. This challenging dataset maps around 1.4k compounds, including small molecules, metals, and salts, to any combination of 27 side effects - leading performers are only slightly better than random guessing (AUROC 0.5).
Among the best performers in literature is the previously discussed CMPNN [8] with a reported AUROC of 0.666+/-0.007, which narrowly outperforms Chemprop at 0.646+/-0.016. Using the same approach, fastprop achieves a decent AUROC of 0.636+/-0.019. Despite many of the entries in this dataset being atypical for mordred near-leading performance is still possible, supporting the robustness and generalizability of this framework.
Pgp
The model in the original study uses a molecular interaction field but has since been surpassed by other models. According to TDC the current leader [38] on this benchmark has achieved an AUROC of 0.938 0.002 555See the TDC Pgp leaderboard.. AOn the same leaderboard Chemprop [6] achieves 0.886 0.016 with the inclusion of additional molecular features. fastprop yet again approaches the performance of the leading methods and outperforms Chemprop, here with an AUROC of 0.919 0.013 and an accuracy of 84.5 0.2%.
ARA
The final benchmark is one which closely mimics typical QSPR studies. Compiled by Schaduangrat et al. in 2023 [37], this dataset maps ~0.8k small molecules to a binary label indicating if the molecule is an Androgen Receptor Antagonist (ARA). The reference study introduced DeepAR, a highly complex modeling approach, which achieved an accuracy of 0.911 and an AUROC of 0.945.
For this study an 80/10/10 random splitting is repeated four times on the dataset since no analogous split to the reference study can be determined. Chemprop takes 16 minutes and 55 seconds to run on this dataset and achieves only 0.824+/-0.020 accuracy and 0.898+/-0.022 AUROC. fastprop takes only 2 minutes and 4 seconds (1 minute and 47 seconds for descriptor calculation) and is competitive with the reference study in performance, achieving a 89.1+/-4.0% accuracy and 0.951+/-0.018 AUROC.
Limitations and Future Work
@@
Although fastprop is consistently around an order of magnitude faster to train than learned representations when using a GPU, execution time is a minor concern when considering the enormous labor invested in dataset generation. For day-to-day work it is convenient but the correctness of fastprop, especially on small datasets, is more important. Note that due to the large size of the FNN in fastprop it will typically be slower than Chemprop when training on a CPU since Chemprop uses a much smaller FNN and associated components.
Regardless, there is an clear performance improvement to be had by reducing the number of descriptors to a subset of only the most important. Future work will address this possibility to decrease time requirements for both training by reducing network size and inference by decreasing the number of descriptors to be calculated for new molecules. This has not been done in this study for two reasons: (1) to emphasize the capacity of the DL framework to effectively perform feature selection on its own via the training process, de-emphasizing unimportant descriptors; (2) as discussed above, training time is small compared ot dataset generation time.
@@
Though not discussed here for the sake of length, fastprop already contains the functionality to perform feature importance studies on trained models. By using SHAP values [45] to assign a scalar ‘importance’ to each of the input features, users can determine which of the mordred descriptors has the largest impact on model predictions. Future studies will demonstrate this in greater detail.
@@ Acknowledgements
pre 28. Saldana DA, Starck L, Mougin P, et al (2011) Flash point and cetane number predictions for fuel compounds using quantitative structure property relationship (QSPR) methods. Energy & Fuels 25:3900–3908. https://doi.org/10.1021/ef200795j
pre 29. Das DD, St. John PC, McEnally CS, et al (2018) Measuring and predicting sooting tendencies of oxygenates, alkanes, alkenes, cycloalkanes, and aromatics on a unified scale. Combustion and Flame 190:349–364. https://doi.org/10.1016/j.combustflame.2017.12.005
pre 30. Eibeck A, Nurkowski D, Menon A, et al (2021) Predicting power conversion efficiency of organic photovoltaics: Models and data analysis. ACS Omega 6:23764–23775. https://doi.org/10.1021/acsomega.1c02156
pre 31. Bhat V, Sornberger P, Pokuri BSS, et al (2023) Electronic, redox, and optical property prediction of organic pi-conjugated molecules through a hierarchy of machine learning approaches. Chemical Science 14:203–213. https://doi.org/10.1039/d2sc04676h
pre 32. Delaney JS (2004) ESOL: Estimating aqueous solubility directly from molecular structure. Journal of Chemical Information and Computer Sciences 44:1000–1005. https://doi.org/10.1021/ci034243x
pre 33. Song Y, Zheng S, Niu Z, et al (2020) CMPNN README
pre 34. Mobley DL, Guthrie JP (2014) FreeSolv: A database of experimental and calculated hydration free energies, with input files. Journal of Computer-Aided Molecular Design 28:711–720. https://doi.org/10.1007/s10822-014-9747-x
pre 35. Lopez SA, Pyzer-Knapp EO, Simm GN, et al (2016) The harvard organic photovoltaic dataset. Scientific Data 3: https://doi.org/10.1038/sdata.2016.86
pre 36. Burns JW, Rogers DM (2023) QuantumScents: Quantum-mechanical properties for 3.5k olfactory molecules. Journal of Chemical Information and Modeling 63:7330–7337. https://doi.org/10.1021/acs.jcim.3c01338
pre 37. Schaduangrat N, Anuwongcharoen N, Charoenkwan P, Shoombuatong W (2023) DeepAR: A novel deep learning-based hybrid framework for the interpretable prediction of androgen receptor antagonists. Journal of Cheminformatics 15: https://doi.org/10.1186/s13321-023-00721-z
pre 38. Notwell JH, Wood MW (2023) ADMET property prediction through combinations of molecular fingerprints
pre 39. Riesen K, Bunke H (2008) IAM graph database repository for graph based pattern recognition and machine learning. In: Vitoria Lobo N da, Kasparis T, Roli F, et al (eds) Structural, syntactic, and statistical pattern recognition. Springer Berlin Heidelberg, Berlin, Heidelberg, pp 287–297
pre 40. Szymański P, Kajdanowicz T (2017) A network perspective on stratification of multi-label data. In: Luís Torgo PB, Moniz N (eds) Proceedings of the first international workshop on learning with imbalanced domains: Theory and applications. PMLR, pp 22–35
pre 41. Kuhn M, Letunic I, Jensen LJ, Bork P (2015) The SIDER database of drugs and side effects. Nucleic Acids Research 44:D1075–D1079. https://doi.org/10.1093/nar/gkv1075
pre 42. Broccatelli F, Carosati E, Neri A, et al (2011) A novel approach for predicting p-glycoprotein (ABCB1) inhibition using molecular interaction fields. Journal of Medicinal Chemistry 54:1740–1751. https://doi.org/10.1021/jm101421d
pre 43. Huang K, Fu T, Gao W, et al (2021) Therapeutics data commons: Machine learning datasets and tasks for therapeutics. CoRR abs/2102.09548: https://doi.org/10.48550/arXiv.2102.09548
pre 44. Carbonell P, Carlsson L, Faulon J-L (2013) Stereo signature molecular descriptor. Journal of Chemical Information and Modeling 53:887–897. https://doi.org/10.1021/ci300584r
pre 45. Lundberg S, Lee S-I (2017) A unified approach to interpreting model predictions

### v2

@@ Introduction
For this unique dataset, execution time trends are inverted. fastprop takes 1 minute and 43 seconds, of which 1 minute and 31 seconds were spent calculating descriptors for these unusually large molecules. Chemprop completes in 1 minute and 16 seconds, faster overall but much slower compared with the training time of fastprop without descriptor calculation.
Limitations and Future Work
Negative Results
Delta Learning with Fubrain
First described by Esaki and coauthors, the Fraction of Unbound Drug in the Brain (Fubrain) dataset is a collection of about 0.3k small molecule drugs and their corresponding experimentally measured unbound fraction in the brain, a critical metric for drug development [22]. This specific target in combination with the small dataset size makes this benchmark highly relevant for typical QSPR studies, particular via delta learning. DeepDelta [14] performed a 90/0/10 cross-validation study of the Fubrain dataset in which the training and testing molecules were intra-combined to generate all possible pairs and then the differences in the property 555Although the original Fubrain study reported untransformed fractions, the DeepDelta authors confirmed via GitHub that DeepDelta was trained on log base-10 transformed fraction values, which is replicated here. were predicted rather than absolute values, increasing the amount of training data by a power of two.
DeepDelta reported an RMSE of 0.830 0.023, whereas a Chemprop model trained to directly predict property values was only able to reach an accuracy of 0.965 0.019 when evaluated on its capacity to predict property differences. fastprop is able to outperform Chemprop, though not DeepDelta, achieving an RMSE of 0.930 0.029 when using the same splitting procedure above. It is evident that delta learning is still a powerful technique for regressing small datasets.
For completeness, the performance of Chemprop and fastprop on Fubrain are also compared to the original study. The study that first generated this dataset used mordred descriptors but as is convention they strictly applied linear modeling methods. Using both cross validation and and external test sets, they had an effective training/validation/testing split of 0.64/0.07/0.28 which will be repeated 4 times here for comparison. All told, their model achieved an RMSE of 0.53 averaged across all testing data. In only 39 seconds, of which 31 are spent calculating descriptors, fastprop far exceeds the reference model with an RMSE of 0.207 0.024. This also surpasses Chemprop, itself outperforming the reference model with an RMSE of 0.223 0.036.
fastprop Fails on QuantumScents
@@
In the reference study, Chemprop achieved an AUROC of 0.85 with modest hyperparameter optimization and an improved AUROC of 0.88 by incorporating the atomic descriptors calculated as part of QuantumScents. fastprop is incapable of incorporating atomic features, so they are not included. Using only the 2D structural information, fastprop falls far behind the reference study with an AUROC of only 0.651 0.005. Even when using the high-quality 3D structures and calculating additional descriptors (demonstrated in the GitHub repository), the performance does not improve.
The exact reason for this failure is unknown. Possible reasons include that the descriptors in mordred are simply not correlated with this target, and thus the model struggles to make predictions. This is a fundamental drawback of this fixed representation method - whereas a LR could adapt to this unique target, fastprop fails.
@@
fastprop is consistently faster to train than Chemprop when using a GPU, helping exploit the ‘time value’ of data. Note that due to the large size of the FNN in fastprop it can be slower than small Chemprop models when training on a CPU since Chemprop uses a much smaller FNN and associated components.
There is a clear performance improvement to be had by reducing the number of descriptors to a subset of only the most important. Future work can address this possibility to decrease time requirements for both training by reducing network size and inference by decreasing the number of descriptors to be calculated for new molecules. This has not been done in this study for two reasons: (1) to emphasize the capacity of the DL framework to effectively perform feature selection on its own via the training process, de-emphasizing unimportant descriptors; (2) as discussed above, training time is small compared ot dataset generation time.
@@
Though not discussed here for the sake of length, fastprop already contains the functionality to perform feature importance studies on trained models. By using SHAP values [37] to assign a scalar ‘importance’ to each of the input features, users can determine which of the mordred descriptors has the largest impact on model predictions. The utility of these values can be explored in greater detail on a case-by-case basis.
@@ Acknowledgements
pre 28. Notwell JH, Wood MW (2023) ADMET property prediction through combinations of molecular fingerprints
pre 29. Schaduangrat N, Anuwongcharoen N, Charoenkwan P, Shoombuatong W (2023) DeepAR: A novel deep learning-based hybrid framework for the interpretable prediction of androgen receptor antagonists. Journal of Cheminformatics 15: https://doi.org/10.1186/s13321-023-00721-z
pre 30. Saldana DA, Starck L, Mougin P, et al (2011) Flash point and cetane number predictions for fuel compounds using quantitative structure property relationship (QSPR) methods. Energy & Fuels 25:3900–3908. https://doi.org/10.1021/ef200795j
pre 31. Das DD, St. John PC, McEnally CS, et al (2018) Measuring and predicting sooting tendencies of oxygenates, alkanes, alkenes, cycloalkanes, and aromatics on a unified scale. Combustion and Flame 190:349–364. https://doi.org/10.1016/j.combustflame.2017.12.005
pre 32. Broccatelli F, Carosati E, Neri A, et al (2011) A novel approach for predicting p-glycoprotein (ABCB1) inhibition using molecular interaction fields. Journal of Medicinal Chemistry 54:1740–1751. https://doi.org/10.1021/jm101421d
pre 33. Huang K, Fu T, Gao W, et al (2021) Therapeutics data commons: Machine learning datasets and tasks for therapeutics. CoRR abs/2102.09548: https://doi.org/10.48550/arXiv.2102.09548
pre 34. Burns JW, Rogers DM (2023) QuantumScents: Quantum-mechanical properties for 3.5k olfactory molecules. Journal of Chemical Information and Modeling 63:7330–7337. https://doi.org/10.1021/acs.jcim.3c01338
pre 35. Szymański P, Kajdanowicz T (2017) A network perspective on stratification of multi-label data. In: Luís Torgo PB, Moniz N (eds) Proceedings of the first international workshop on learning with imbalanced domains: Theory and applications. PMLR, pp 22–35
pre 36. Carbonell P, Carlsson L, Faulon J-L (2013) Stereo signature molecular descriptor. Journal of Chemical Information and Modeling 53:887–897. https://doi.org/10.1021/ci300584r
pre 37. Lundberg S, Lee S-I (2017) A unified approach to interpreting model predictions